# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ Appendix A Proofs of Lemmas
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, then $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$. Q.E.D.

### v3

@@ Appendix A Proofs of Lemmas
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{j}=-\sinh q_{k}\cdot\sinh\theta_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$, and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, it follows that $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$. Q.E.D.